# An Empirical Characterization of Outages and Incidents in Public Services for Large Language Models

### v1

@@ Abstract.
Failure characterization, LLM, performance modeling, reliability, OpenAI, Anthropic, Character.AI, operational data analytics, outage, incident, failure-recovery, time-series analysis
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXcopyright: noneconference: 16th ACM/SPEC International Conference on Performance Engineering; May 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
We identify and address in this work two main challenges in understanding how public LLM services currently fail. First, no longitudinal service failure data currently exists. Ideally, the community would have access to a large number of similarly curated datasets that capture LLM-service failures, under the same failure model, over long periods of time. There are some efforts to provide available LLM workloads, such as BurstGPT (Wang et al., 2024a) and AcmeTrace (Hu et al., 2024), but they each focus on one LLM service, and none provides service failure data for it.
Second, no comprehensive analysis of failures in public LLM services currently exists. At this stage in the scientific area, such an analysis would ideally be data-driven, and include for example general characteristics of failures, such as Mean Time Between Failures (MTBF) and To Recovery (MTTR) from classical dependability analysis (Avizienis et al., 2004), and also of the time spent in various stages of the recovery-process specific to LLM operations; a temporal analysis of failures; and an analysis of failure cascades (co-occurrences). Such kinds of analysis would enable future research into models, and future theoretical and practical studies of LLM systems.
@@
    We summarize the de facto industry standard for modeling LLM-service outages and exemplify the anatomy of an outage (Section 2).
@@
    We collect outage and incident data for 8  LLM services, and prepare the corresponding LLM-failure datasets (Section 3). This study covers representative, commonly used LLM services, across 3 LLM-service providers: OpenAI’s API (API, 2024b), ChatGPT (ChatGPT, 2024), DALL·E (DALL·E, 2024), Playground (Playground, 2024); Anthropic’s API (API, 2024a), Claude (Claude, 2024), Console (Console, 2024); and Character.AI (Character.AI, 2024).
@@
    We analyze the failure characteristics of 8 LLM services (Section 4). We analyze the MTTR and MTBF by provider and by service, the time spent in various stages of the recovery process, and quantify empirically the model parameters introduced in Section 2.
@@
    We analyze LLM service failures over time (Section 5). We explore service availability over hourly and daily intervals, identify various diurnal and weekly patterns, and investigate auto-correlations.
@@
    We analyze the co-occurrence of failures (Section 6). Specifically, we analyze the co-occurrence of failures per provider, and of pairs of services across providers.
@@
    We follow the principles of open science and release as open, FAIR artifacts for public use the datasets [^1] https://zenodo.org/records/14018219 and software [^2] https://github.com/atlarge-research/llm-service-analysis used in this work. We expect them to provide opportunities to reproduce and expand on our research.
@@ 2.1. Model and Real-World Example
As Section 2.2 details, the model also includes LLM-specific metrics to capture the outage duration, scale it according to the severity of the outage, and estimate the (daily) service availability.
@@ 3.2. Data Collection and Dataset Preparation
We collected all available outage and incident data reported publicly by of OpenAI, Anthropic, and Character.AI, up to 2024-08-31, on their public status pages (Uptime, 2024c, a, b) and incident pages (Incident, 2024c, a, b). The starting dates differ: OpenAI has started reporting on February 11, Anthropic on July 11, and Character.AI on October 19, all dates in 2023. (Our study misses none of the published reports.)
@@
We developed an automated data-collection method, able to collect industry-standard outage and incident reports. Our tools leverage Python Selenium WebDriver (Selenium, 2024), a robust tool allowing native browser automation by simulating real-user interactions. Our tools implement exception-handling mechanisms, addressing potential issues such as network problems, stale elements, and unexpected page layouts. They parse and extract information from the dynamic pages, and store them as raw datasets.
To prepare the datasets, we did typical data transformation, including filling in missing values, extracting data from text, processing JSON formats, splitting columns, and feature engineering to get the metrics used in this study. Last, before the analysis, we performed data cleanup, as discussed in the following, to obtain the cleaned outage and incident datasets whose properties are summarized in Table 3 and Table 4, respectively.
@@ 4. Failure-Recovery Analysis
Observation #3: Claude spent the longest time on the periods of investigating (1.35 hours), repairing (1.72 hours), and checking (2.05 hours).
Figure 4 shows the percent of the 3 periods (Investigating, Repairing, Checking) in $MTTR$. The majority of the resolution time is used for checking, ranging from the highest 82.71% for Character.AI to the lowest 43.75% for Console. Most services spent more percent of the time on repairing than investigating, except for API and Console from Anthropic. Anthropic API spent 26.70% of the time investigating issues, while Console spent 43.23% in the same period.
The large fraction of the time used for checking indicates that deploying a fix to production takes a long time. Operators should employ faster testing and continuous deployment techniques to deploy fixes faster (Zhang et al., 2018). However, this is challenging as LLMs are a new technology, and there isn’t much work on improving testing and deployment time.
@@
Although Anthropic resolves failures the fastest, it also encounters them most frequently, with every 5.22 days on average. In contrast, OpenAI and Character.AI are more reliable, with failure occurring every 8.48 and 8.74 days, respectively. A notable percentage of incidents occur within a day: 35.47% for Anthropic, 28.77% for OpenAI, and 20.00% for Character.AI. Within 1 week interval, nearly three-quarters of failures occur for OpenAI (75.64%) and for Anthorpic (78.63%), with a slightly lower rate for Character.AI (60.00%). Over 90% of incidents for all providers happen within a month of each other.
Our findings indicate that users of LLMs should expect failures regularly (at least once a month). Therefore, failure should not be an exceptional event but should be baked into the users’ normal operating procedure.
@@ 5. Failure Patterns Over Time
5.1. Temporal Distributions
Observation #8: OpenAI and Anthropic’s services failures exhibit periodic patterns that more frequent on weekdays than on weekends. However, Character.AI has fewer failures on Tuesdays and Wednesdays. All services show a diurnal pattern of failures, typically peaking from 8:00 to 16:00.
@@ 5.2. Auto-correlations
We investigate if a failure is immediately followed by another failure and how often it happens. Figure 8 depicts the auto-correlation for the number of incidents at month, week, and day granularities. Confidence intervals are drawn as the blue area. By default, this is set to a 95% confidence interval, suggesting that correlation values outside of this area are significant, which are real patterns rather than random noise. Lags represent the time intervals at which a time series is compared to itself, and autocorrelation measures how similar a time series is to itself at different lags.
@@ 6. Co-occurrence of Failures
This section examines the co-occurrence of failures across services. When an outage occurs in one service, do other services also experience outages? How about the impacted range of incidents for different services and providers? To address these questions, we analyze (1) the co-occurrence of outages, and (2) the impact range of incidents.
@@ 6.1. Co-occurrence of Outages
The heatmaps show that co-occurrence is notably high among services from the same provider. For OpenAI services, the API is more likely to have an outage with DALL·E (78.26%) and Playground (87.50%) than ChatGPT (49.21%). For Anthropic’s services, the likelihood of any two services experiencing outages on the same day is extremely high over 80%, this may be caused by a lack of isolation across different services. There is no correlation observed between services from different providers. The lack of correlation suggests that user can use one service as the other’s backup to increase their reliability. The difference in co-occurrence between OpenAI and Anthropic suggests that outages could be reduced through better service isolation.
@@ 9. Conclusion
Understanding the characteristics of failures in the operation of public LLM services has become a stringent problem, driven by the rapid increase in the popularity of such services, market competitiveness, and increasingly self-reported presence of such failures by LLM service providers. Addressing this problem, in this work we have conducted a comprehensive empirical characterization of outages and incidents in public LLM services.
We have collected long-term failure data from 8 relevant services, and produced corresponding outage and incident datasets under the same failure model. We analyzed the failure characteristics of these services, per service and overall, and specifically analyzed MTTR and MTBF, time spent in various stages in the recovery process, service availability over hourly and daily intervals, diurnal and weekly availability distributions, auto-correlations; and failure co-occurrence for pairs and groups of services per provider, and for pairs of services across providers.
In our analysis, we emphasized over 10 observations, which scientists, engineers, and users could include directly in their knowledge base, and from which improvements to LLM systems could occur in time.
For the future, we aim to lead a community effort where datasets, collected long-term and processed to provide similar information, can be shared. Future analysis could include the popular LLM services powered by Google DeepMind’s Gemini and Mistral AI’s Mixtral, and the promising emerging LLMs from US-, China-, and Japan-based big tech companies, e.g., NVIDIA’s Nemotron.
@@ Acknowledgements.
We thank the China Scholarship Council (CSC) for supporting Xiaoyu Chu. We thank the support of Netherlands-funded projects NWO OffSense and GFP 6G FNS, and EU-funded projects MCSA-RISE Cloudstars and Horizon Graph-Massivizer.

### v2

@@ Abstract.
Failure characterization, LLM, failure-recovery, reliability, OpenAI, Anthropic, Character.AI, operational data analytics
copyright: ccjournalyear: 2025copyright: ccconference: Proceedings of the 16th ACM/SPEC International Conference on Performance Engineering; May 5–9, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 16th ACM/SPEC International Conference on Performance Engineering (ICPE ’25), May 5–9, 2025, Toronto, ON, Canadadoi: 10.1145/3676151.3719372isbn: 979-8-4007-1073-5/2025/05ccs: Computer systems organization Reliability
@@ 1. Introduction
We identify and address in this work two main challenges in understanding how public LLM services currently fail. First, no longitudinal service failure data currently exists. Ideally, the community would have access to a large number of similarly curated datasets that capture LLM-service failures, under the same failure model, over long periods of time. There are some efforts to provide available LLM workloads, such as BurstGPT (Wang et al., 2024a) and AcmeTrace (Hu et al., 2024), but they each focus on one LLM service, and none provides service failure data for it. Second, no comprehensive analysis of failures in public LLM services currently exists. At this stage in the scientific area, such an analysis would ideally be data-driven, and include for example general characteristics of failures, such as Mean Time Between Failures (MTBF) and To Recovery (MTTR) from classical dependability analysis (Avizienis et al., 2004), and also of the time spent in various stages of the recovery-process specific to LLM operations; a temporal analysis of failures; and an analysis of failure cascades (co-occurrences). Such kinds of analysis would enable future research into models, and future theoretical and practical studies of LLM systems.
@@
    We summarize the de facto industry standard for modeling LLM-service outages and exemplify the anatomy of an outage (Section 2);
@@
    We collect outage and incident data for 8  LLM services, and prepare the corresponding LLM-failure datasets (Section 3). This study covers representative, commonly used LLM services, across 3 LLM-service providers;
@@
    We analyze the failure characteristics of 8 LLM services (Section 4). We analyze the MTTR and MTBF by provider and by service, the time spent in various stages of the recovery process, and quantify empirically the model parameters;
@@
    We analyze LLM service failures over time (Section 5). We explore service availability over hourly and daily intervals, identify various diurnal and weekly patterns, and investigate auto-correlations;
@@
    We analyze the co-occurrence of failures (Section 6). Specifically, we analyze the co-occurrence of failures per provider, and of pairs of services across providers;
@@
    We follow the principles of open science and release the datasets and software as open, FAIR artifacts to enable reproducibility and further research.
@@ 3.2. Data Collection and Dataset Preparation
We collected all available outage and incident data reported publicly by of OpenAI, Anthropic, and Character.AI, up to 2024-08-31, on their public status pages (Uptime, 2024c, a, b) and incident pages (Incident, 2024c, a, b). The starting dates differ: OpenAI has started reporting on February 11, Anthropic on July 11, and Character.AI on October 19, all dates in 2023. Our study misses none of the published reports.
@@
We developed an automated data-collection method, able to collect industry-standard outage and incident reports. Our tools leverage Python Selenium WebDriver (Selenium, 2024), a robust tool allowing native browser automation by simulating real-user interactions. Our tools implement exception-handling mechanisms, addressing potential issues such as network problems, stale elements, and unexpected page layouts. They parse and extract information from the dynamic pages and store them as raw outage and incident datasets. After that, we performed a series of data transformations for the raw datasets, including filling in missing values, extracting data from text, processing JSON formats, splitting columns, and performing feature calculations to get the metrics used in this study.
@@ 4. Failure-Recovery Analysis
Observation #3: Claude spent the longest time on investigating (1.35 hours), repairing (1.72 hours), and checking (2.05 hours).
Figure 4 shows the percent of the 3 periods (Investigating, Repairing, Checking) in $MTTR$. The majority of the resolution time is used for checking, ranging from the highest 82.71% for Character.AI to the lowest 43.75% for Console. Most services spent more percent of the time on repairing than investigating, except for API and Console from Anthropic. Anthropic API spent 26.70% of the time investigating issues, while Console spent 43.23% in the same period. The large fraction of the time used for checking indicates that deploying a fix to production takes a long time. Operators should employ faster testing and continuous deployment techniques to deploy fixes faster (Zhang et al., 2018). However, this is challenging as LLMs are a new technology, and there isn’t much work on improving testing and deployment time.
@@
Although Anthropic resolves failures the fastest, it also encounters them most frequently, with every 5.22 days on average. In contrast, OpenAI and Character.AI are more reliable, with failure occurring every 8.48 and 8.74 days, respectively. A notable percentage of incidents occur within a day: 35.47% for Anthropic, 28.77% for OpenAI, and 20.00% for Character.AI. Within 1 week interval, nearly three-quarters of failures occur for OpenAI (75.64%) and for Anthorpic (78.63%), with a slightly lower rate for Character.AI (60.00%). Over 90% of incidents for all providers happen within a month of each other. Our findings indicate that users of LLMs should expect failures regularly (at least once a month). Therefore, failure should not be an exceptional event but should be baked into the users’ normal operating procedure.
@@ 5. Failure Patterns Over Time
5.1. Temporal Distributions
Observation #8: OpenAI and Anthropic exhibit more failures on weekdays, while Character.AI has fewer on Tuesdays and Wednesdays. All services show peak failures from 8:00 to 16:00.
@@ 5.2. Auto-correlations
We investigate if a failure is immediately followed by another failure and how often it happens. Figure 8 depicts the auto-correlation for the number of incidents at month, week, and day granularities. We use the autocorrelation function (ACF) to measure the degree of correlation based on the temporal incidents data. Confidence intervals are drawn as the blue area. By default, this is set to a 95% confidence interval, suggesting that correlation values outside of this area are significant, which are real patterns rather than random noise. Lags represent the time intervals at which a time series is compared to itself, and autocorrelation measures how similar a time series is to itself at different lags.
@@ 6. Co-occurrence of Failures
This section examines the co-occurrence of failures across services. When an outage occurs in one service, do other services also experience outages? Is there any co-occurrence within and across different providers, given that services may share the same cloud infrastructure? For instance, both Anthropic and Character.AI rely on GCP (gcp, 2025b, a). How about the impacted range of incidents for different services and providers? To address these questions, we analyze (1) the co-occurrence of outages, and (2) the impact range of incidents.
@@ 6.1. Co-occurrence of Outages
The heatmaps show that co-occurrence is notably high among services from the same provider. For OpenAI services, the API is more likely to have an outage with DALL·E (78.26%) and Playground (87.50%) than ChatGPT (49.21%). For Anthropic’s services, the likelihood of any two services experiencing outages on the same day is extremely high over 80%, this may be caused by a lack of isolation across different services. There is no correlation observed between services from different providers. The lack of correlation suggests that user can use one service as the other’s backup to increase their reliability. The difference in co-occurrence between OpenAI and Anthropic suggests that outages could be reduced through better service isolation. Based on publicly available information, this may be caused by the different cloud infrastructures used by LLMs: OpenAI relies on Azure (azu, 2025), while Anthropic uses GCP (gcp, 2025b).
@@ 9. Conclusion
Understanding the characteristics of failures in the operation of public LLM services has become a stringent problem, driven by the rapid increase in the popularity of such services, market competitiveness, and increasingly self-reported presence of such failures by LLM service providers. Addressing this problem, in this work we have conducted a comprehensive empirical characterization of long-term outages and incidents in public LLM services.
Our main findings includes: (1) Different LLM services take varying amounts of time at different stages of the failure-recovery process. For example, Anthropic services spent more time for investigating and resolving issues than OpenAI services. (2) OpenAI and Anthropic’s services failures exhibit periodic patterns that are more frequent on weekdays than on weekends. However, Character.AI has fewer failures on Tuesdays and Wednesdays. (3) Co-occurrence is significantly higher among services within the same provider, with no clear co-occurrence observed between services from different providers.
Overall, we emphasized over 10 observations, which scientists, engineers, and users could include directly in their knowledge base, and from which improvements to LLM systems could occur in time. For the future work, we aim to lead a community effort where LLM service availability datasets, collected long-term and processed to provide similar information, can be shared. Future analysis could include promising emerging LLM services in different countries such as Google Gemini, Mistral AI, and DeepSeek.
@@ Acknowledgements.
This work was supported by the EU Horizon Graph Massivizer and the EU MSCA Cloudstars projects. This research was partly supported by a National Growth Fund through the Dutch 6G flagship project ”Future Network Services”. We thank the China Scholarship Council (CSC) for supporting Xiaoyu Chu. We thank Daniel Hofstätter, Krijn Doekemeijer, and Radu Nicolae for their critical review of the manuscript.